# A Survey of the Holroyd-Talbot Conjecture

### v1

@@ 4 HK graphs
As in the case of isolated vertices, it can be part of a useful technique to know where the maximum $r$-centers are located in a graph, especially within inductive or injective methods. In studying the EKR property for connected graphs, it is reasonable to start with the fewest possible edges; i.e. trees. Theorems 27, 28, and 29 are important results in this direction. As has been noted regarding the above results, among known witnesses are minimum degree vertices. It makes sense, therefore, to guess that every tree has some leaf as a maximum center, as was conjectured in [47], and verified for small $r$.
@@
If $T$ is a tree and $r\leq 4$ then some maximum $r$-center of $T$ is a leaf.
However, numerous counterexamples to the conjecture were soon found for $r\geq 5$ [2, 12, 28]. The maximum star centers in [2, 12] were of degree 2. However, Feghali, et al. show that $r$-centers can be far from leaves and have any degree.
@@
Let $S^{t,\ell}$ be the $\ell$-uniform spider with degree-$t$ split vertex $w$. Then define the tree $T=T^{d,t,\ell}$ by joining a vertex $v$ to the split vertices of $d$ copies of $S^{t,\ell-1}$. Through very careful analysis, they are able to show that $|{\cal I}_{v}(T)|>|{\cal I}_{u}(T)|$ for every leaf $u$ when $t$ is large enough. It follows, then, that $|{\cal I}_{v}^{r}(T)|>|{\cal I}_{u}^{r}(T)|$ for some value of $r$. They complete the proof by showing that, for every remaining vertex $x$, $|{\cal I}_{v}^{r}(T)|\geq|{\cal I}_{x}^{r}(T)|$ for all $r$.
@@
If every non-pendant vertex of a graph $G$ has an escape path then $G$ is HK.
@@
If $D$ is a thread dominating set of $G$ then any short $D$-spidering of $G$ is HK.
@@
The authors of [49] conjectured that trees with no degree-2 vertices are HK. As a test case, we find in [32] the following family of graphs suggested. For $1\leq i\leq 3$, let $T_{i}(h)$ be a complete binary tree of depth $h$, having unique degree-2 vertex $v_{i}$. Now join a new vertex $v$ to each $v_{i}$ to create the tree $T(h)$, in which every interior vertex has degree 3.
@@
They also construct a caterpillar $P^{k,t}$, generalizing $P_{k}^{*}=P^{k,1}$ by attaching $t$ pendant vertices to each vertex of $P_{k}=v_{1}\cdots v_{k}$. They prove, for all $r$ and $t$, that every pendant neighbor of $v_{2}$ or $v_{k-1}$ is a maximum $r$-center (the $t=1$ case had been done in [19]).
@@ 5 Chvátal’s Conjecture
It is a monument to the difficulty of the conjecture that it is so well-known and yet so few papers have addressed it in the half century since its birth. There were some early papers. Kleitman and Magnanti [52] solved the case in which ${\cal F}$ is the union of two stars. Sterboul solved the rank at most 3 case, where the rank of ${\cal F}$ is the size of its largest set. In 1976, Schönheim [61] solved case in which the maximal elements of ${\cal F}$ intersect. Then there was a long gap until Snevily [63] solved the case in which ${\cal F}$ is 1-shifted in 1992, and quite another long period before Borg [10] generalized the conjecture to weighted families in 2011, solving the 1-shifted case as well. Then in 2017, the paper [17] reproved Sterboul’s result, once with a relatively similar approach, and once with the Sunflower Lemma, in hopes that the method might yield more fruit. Subsequently, Olarte, Santos, and Spreer [60] reproved both Kleitman-Magnanti and Sterboul with shorter proofs. Most recently, Frankl and Kupavskii [33] solved the case in which ${\cal F}$ has cover number 2.
@@ 6 Questions and Remarks
      What Cartesian (or other) products of various graphs are $r$-EKR; e.g., products of paths, cycles, or complete graphs? For some graph product $\times$, could it be that, if $G$ $H$ are $r$-EKR for all $r$, then is $G\times H$ $r$-EKR for all $r$?
@@
      [25] Where are the maximum $r$-centers for caterpillars?

### v3

@@ 4 HK graphs
As in the case of isolated vertices, it can be part of a useful technique to know where the maximum $r$-centers are located in a graph, especially within inductive or injective methods. In studying the EKR property for connected graphs, it is reasonable to start with the fewest possible edges; i.e. trees. Theorems 29, 30, and 31 are important results in this direction. As has been noted regarding the above results, among known witnesses are minimum degree vertices. It makes sense, therefore, to guess that every tree has some leaf as a maximum center, as was conjectured in [47], and verified for small $r$.
@@
If $T$ is a tree and $r\leq 4$, then some maximum $r$-center of $T$ is a leaf.
However, numerous counterexamples to the conjecture were soon found for $r\geq 5$ [2, 12, 28]. The maximum star centers in [2, 12] were of degree 2. However, Feghali, Johnson, and Thomas [28] showed that $r$-centers can be far from leaves and have any degree.
@@
Let $S^{t,\ell}$ be the $\ell$-uniform spider with split vertex of degree $t$. Then define the tree $T=T^{d,t,\ell}$ by joining a vertex $v$ to the split vertices of $d$ copies of $S^{t,\ell-1}$. Through very careful analysis, they are able to show that $|{\cal I}_{v}(T)|>|{\cal I}_{u}(T)|$ for every leaf $u$ when $t$ is large enough. It follows, then, that $|{\cal I}_{v}^{r}(T)|>|{\cal I}_{u}^{r}(T)|$ for some value of $r$. They complete the proof by showing that, for every remaining vertex $x$, $|{\cal I}_{v}^{r}(T)|\geq|{\cal I}_{x}^{r}(T)|$ for all $r$.
@@
If every non-pendant vertex of a graph $G$ has an escape path, then $G$ is HK.
@@
If $D$ is a thread dominating set of $G$, then any short $D$-spidering of $G$ is HK.
@@
The authors of [49] conjectured that trees with no degree-2 vertices are HK. As a test case, the authors of [32] propose showing that the following family of graphs is $r$-EKR. For $1\leq i\leq 3$, let $T_{i}(h)$ be a complete binary tree of depth $h$, having unique degree-2 vertex $v_{i}$. Now join a new vertex $v$ to each $v_{i}$ to create the tree $T(h)$, in which every interior vertex has degree 3.
@@
They also construct a caterpillar $P^{k,t}$ by attaching $t$ pendant vertices to each vertex of $P_{k}=v_{1}\cdots v_{k}$; thus $P^{k,1}=P_{k}^{*}$, the fully pendant contruction for well-covered graphs. They prove, for all $r$ and $t$, that every pendant neighbor of $v_{2}$ or $v_{k-1}$ is a maximum $r$-center (the $t=1$ case had been done in [19]).
@@ 5 Chvátal’s Conjecture
It is a monument to the difficulty of the conjecture that it is so well-known and yet so few papers have addressed it in the half century since its birth. There were some early papers. Kleitman and Magnanti [53] solved the case in which ${\cal F}$ is the union of two stars. Sterboul [66] solved the rank at most 3 case, where the rank of ${\cal F}$ is the size of its largest set. In 1976, Schönheim [62] solved case in which the maximal elements of ${\cal F}$ intersect. Then there was a long gap until Snevily [64] solved the case in which ${\cal F}$ is 1-shifted in 1992, and quite another long period before Borg [10] generalized the conjecture to weighted families in 2011, solving the 1-shifted case as well. Then in 2017, the paper [17] reproved Sterboul’s result, once with a relatively similar approach, and once with the Sunflower Lemma, in hopes that the method might yield more fruit. Subsequently, Olarte, Santos, and Spreer [61] reproved both Kleitman-Magnanti and Sterboul with shorter proofs. Most recently, Frankl and Kupavskii [33] solved the case in which ${\cal F}$ has cover number 2.
@@ 6 Questions and Remarks
      What Cartesian (or other) products of various graphs are $r$-EKR; e.g., products of paths, cycles, or complete graphs? For some graph product $\times$, could it be that, if $G$ and $H$ are $r$-EKR for all $r$, then is $G\times H$ $r$-EKR for all $r$?
@@
      Where are the maximum $r$-centers for caterpillars? (See [25].)